### v1

# PowerPlots: An Open Source Power Grid Visualization and Data Analysis Framework for Academic Research

Abstract
Data visualization is important for developing an understanding of a complex system. PowerPlots.jl is a data visualization tool for power grids, one of the most complex systems in the world. The design of PowerPlots.jl is intended to facilitate exploration of power grid data while performing research and to facilitate communication of research findings to an audience. Several tools created to support this software also facilitate analysis of power grid data by transforming the data into graph topology or data-frame data formats that are more compatible for some applications. The high level of flexibility in PowerPlots.jl enables researchers who are developing and analyzing methods for solving novel power grid problems to better understand and communicate the complexities of their research.
I Introduction
The power grid is a complex engineered network system and its operation is governed by the physics of power flow. Developing an intuition for the operation of the power grid is a valuable skill that allows a researcher to identify when a unique behavior is occurring or more easily identify an error in their computational method. Power grid visualization allows the researcher to see and recognize these patterns [1].
Visualization of the power grid has two primary values. 1) Exploration: It allows for exploration of data and debugging of methods by quickly and interactively visualizing data, 2) Communication: it enables sharing of specific information to communicate research findings. The two goals of PowerPlots.jl are to enable the exploration of a research problem and the communication of research results.
Many software packages, libraries, or ecosystems exist that can also perform power grid visualization, but none cohesively possess all of the additional flexible capabilities needed for supporting research applications, such as the ability to view and explore custom data, extend data formats with new information, and extensively customize figures for publication. PyPSA [2] or Python for Power System Analysis is an open source energy system modeling project. It supports a range of standard power grid optimization models like unit commitment and capacity expansion, and includes cross sector energy systems modeling. It includes plotting of networks for its users to visualize the results and view the solutions of these models. OpenDSS [3] is a distribution network simulator developed by EPRI. Originally a scripting tool to define networks and routines for simulation, an extension for a graphical environment was developed, OpenDSS-G, to enable a visual interface to create and view simulation results. PowerGridModel [4] is a an open source power grid analysis library that supports power flow, state estimation, and short circuit analysis. It includes visualization software in its data analysis extension [5]. PandaPower [6] is an open source python package for power system analysis. It emphasizes a tabular data structure for data analysis of power flow, state estimation, and short-circuit calculations. It has basic support for visualizing the network representation of the data.
PowerWorld [7] is an industry software initially developed to provide students with a visual tool to understand the power system concepts that they learned in classes. The slogan “the visual approach to electric power systems” emphasizes its roots as a tool to develop intuition outside of looking at charts and tables of power grid data. A variety of other commercial industry software including PSSE [8], PLSF [9], ETAP [10], and TARA [11] have some level of visualization of network data like prices or outage reliability. These tools typically have very limited flexibility to visualize data beyond what they are designed to calculate.
These existing visualization software systems represent a mixture of open and closed source tools. They provide visualization for projects within a specific ecosystem and are generally for visualizing traditional well-studied power grid problems. For users of these ecosystems who are analyzing traditional problems, these are excellent tools. However, they are not always suitable or adaptable for research applications that solve novel power grid problems.
PowerPlots.jl is a flexible tool for creating visualizations of power grids to better support the needs of researchers. It has a simple interface for creating a network plot to view and interact with data, and a high level of customization to create publication-ready figures. It is written in the Julia programming language [12], and adopts the data format from PowerModels.jl [13] and PowerModelsDistribution.jl [14], which parse file formats including MatPower [15], PSSE [8], as well as a JSON dictionary. The data format is extendable to visualize new components that are not natively supported by these file formats. In addition, several tools like a graph data model and a data-frame data model, created to support the network plots, are also useful as data analysis tools for research.
Many grid research packages can take advantage of PowerPlots.jl to enable analysis of their outputs with no change to their code [16, 17, 18, 19], and many publications already use PowerPlots.jl to communicate their results in application areas ranging from wildfire risk, to carbon intensity metrics, restoration planning, and network reconfiguration  [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
This paper is not documentation for PowerPlots.jl; the documentation can accessed at the GitHub repository of the package [^1] https://wispo-pop.github.io/PowerPlots.jl/dev/. Rather, this paper explores how the package can be used as a research tool, including examples of the types of analyses that this package enables.
The primary contributions of this paper are (1) An introduction to the plotting tool PowerPlots.jl and its software design, (2) Examples of how to use PowerPlots.jl to explore network data and customize figures, and (3) An exploration of the types of data analysis that are provided by tools within PowerPlots.jl.
The rest of the paper is organized as follows: Section II explains the design and basic software structure of PowerPlots.jl. Section III describes the features of PowerPlots.jl to enable data exploration, with examples of how to investigate different types of research questions. Section III-F demonstrates the customization that can be achieved using PowerPlots.jl. Section IV shows how the data structures within PowerPlots.jl can be utilized to conduct data analyses of power grid data. Finally, Section V concludes the work.
II Design of PowerPlots
The two goals of PowerPlots.jl are to provide easy exploration of power grid data, and to create clear communication of research results for publication and presentation. To achieve these goals, the design of PowerPlots.jl revolves around these principles: 1) simplicity, 2) flexibility, and 3) customization.
The primary user of PowerPlots.jl is a researcher exploring novel research problems, and these principles enable them to conduct their research more effectively. PowerPlots.jl must be very simple to use with intuitive default behavior for initial data exploration about the network. Flexibility in visualizations allows the user to visualize a wide range of custom information about a power grid. High levels of customization allows a user to make clear visualizations that highlight specific data for an audience.
This section introduces how the software package’s functions to support these principles. First, we introduce an overview of the package functionality. Next, we introduce the data models used by PowerPlots.jl that also have utility for other analytical applications.
II-A Plotting Process
PowerPlots.jl uses the nested dictionary structure of PowerModels.jl as input. The top level of the dictionary contains metadata about the power grid and a key for each of the power grid components. Each grid component key refers to a dictionary containing data for each component.
grid_data = Dict(
bus => Dict(
”1” => Dict(
vmax => 1.1,
vmin => 0.9,
base_kv => 220,
”2” => Dict( )
gen => Dict( )
[…]
The default use of power plots automatically applies these steps to create a figure. However, each step can be customized to enable control over the graph layout, the set of components included in the figure, and which data fields are used for the visualization. If a user requires additional customization beyond what can be achieved with the user arguments, they can then modify the figure to fully customize the representation of e.g legend placement, add data fields to visualize, or change the shape of grid components. As a VegaLite.jl figure, a wide variety of customization can be applied.
II-B Data Structures
[…]
II-B1 PowerModelsGraph
The graph structure created for PowerPlots is an undirected graph of all of the edges and vertices seen in a visualization. By default, network components including buses, generators, loads, and shunts are all nodes in the graph structure, and branches, dc lines, switches, and transformers are included as edges. Nodes that are not buses are considered ’connected components’ because they are directly connected to a bus, and an additional ’connector’ edge type is added to represent these connection points.
The data structure contains four fields: the undirected graph, as well as three mappings from the graph indices to the original power grid data.
mutable struct PowerModelsGraph
graph::Graphs.SimpleDiGraph
node_comp_map::Dict{Int,Tuple{String, String}}
edge_comp_map::Dict{Graphs.AbstractEdge, Tuple{String, String}}
edge_connector_map::Dict{ Graphs.AbstractEdge, Tuple{String, String}}
A user can specify the components to include in the graph. This is useful for computing a network analysis without including information about generator connections. The following example shows how to identify the highest node degree bus in a network by creating a graph where only buses are included as nodes.
using PowerPlots, PGLib, Graphs
grid_data = pglib(”case1354”)
[…]
(v, id) = findmax(degree(pmg.graph)) # (14, 2)
pmg.node_comp_map[id] # (:bus, ”1001”)
# the highest node degree is 14 at bus 1001
II-B2 PowerModelsDataFrame
The data structure used in PowerModels has a nested dictionary structure, with component types at the top level, followed by component indices, with component parameters in a third dictionary. This structure is convenient for many of the common data look-ups and manipulations that are used when exploring power grid data associated with a single component. However, it is not suitable to exploring aggregate data across a component type.
[…]
mutable struct PowerModelsDataFrame
metadata::DataFrames.DataFrame
components::Dict{Symbol, DataFrames.DataFrame}
This data-frame structure allows for easy querying of metrics about the components in the network. The component dictionaries can be accessed and viewed as follows.
using PowerPlots, PGLib
grid_data = pglib(”case1354”)
pmd = PowerModelsDataFrame(grid_data);
pmd.components[:gen]
260×18 DataFrame
Row pg qg gen_bus pmax
1 6.66 1.33 124 10.0
2 6.66 1.35 2035 10.0
3 0.6 0.17 3390 1.2
4 0.4 0.12 1604 0.8
5 13.3 3.08 2446 20.0
6 0.8 0.22 5664 1.6
7 13.3 3.07 5481 20.0
III Exploration of Grid Data
The following section demonstrates several features of PowerPlots.jl that enable exploration of grid data. Each of the examples uses a network from PGLib [31] unless otherwise specified.
Figure 1: 39 Node EPRI Network
III-A Simplicity
First, we demonstrate the simplicity of creating a power grid visualization. The following code example produces the visualization of the EPRI 39 node network in Fig. 1.
using PowerPlots, PGLib
grid_data = pglib(”case39_epri”)
powerplot(grid_data)
This automatically includes the generator, load, branch, and bus components in the network and creates a network layout for visualization. The default settings provide a view of the structure of the network, without visualizing any specific data about the components.
((a)) 89 Node Pegase Network
((b)) 118 Node IEEE Network
Figure 2: It is quickly apparent that the 89 node network in Fig. 2(a) has a tight clustering of nodes in the center of the graph that are densely connected. The 118 node IEEE network is shown in Fig. 2(b) as a comparison of a more typical network that lacks this density.
((a)) Kamada Kawai
((b)) Spring
((c)) SFDP
((d)) Spectral
Two larger networks are shown in Fig. 2, where the differences in the network structure are visually apparent. In Fig. 2(a), there is a very high-density collection of nodes in the center of the graph with a high degree of connectivity between them. In comparison, Fig. 2(b) shows a more typical transmission power grid network that does not have a high degree of connectivity in a subset of the network. This visual observation matches with the network analysis from [32]. In this work, the authors analyze metrics like mean nodal degree, clique size, and adjacency spectra radius, and they estimate that some network processing such as Kron Reduction was applied to the PEGASE 89 node network along with several other PEGASE networks. The same conclusion is evident from viewing the network topology.
Both of these networks are visually very busy with all of the generators, shunts, and loads included in the visualization, and a figure that is used in a paper should be more selective to highlight the network clustering. However, this example demonstrates the default view when running the
powerplot(grid_data)} command without additional arguments.
$\begin$
figure*
Layouts of a 39 node network, sorted by time to compute the layout. The Kamada Kawai layout method generally produces the best visual layout to view the components of the network, but has a much longer computation time. Buses are shown in blue, generators in orange, and loads in red. Lines are shown in green, and dashed gray lines indicate where generators and loads connect to buses.
III-B Layouts
Many synthetic power grid networks used in research do not have labeled coordinates for the components of the power grid. When the coordinates are missing, a network layout must be computed to visualize the power grid. PowerPlots.jl allows several types of network layout algorithms to be used, with varying tradeoffs of visual quality and computational speed. Fig. III-A shows the layout algorithms available in PowerPlots.jl. The default is the Kamada-Kawai algorithm which solves for the geometric distance between nodes by minimizing the deviation from the graph distance between each node pair in the graph [33]. This method produces layouts where each component of the power grid is clearly visible and typically has few edge crossings, however, it is the most computationally expensive of the available layout methods. The other layout methods are provided by NetworkLayout.jl [34]. These methods include the Spring [35], Scalable Force Directed Placement (SFDP) [36], Spectral [37], Shell, and Grid layouts. Each of these methods has parameters, such as a spring force or edge weights, that can modify the resulting layout as seen below.
using PowerPlots, PGLib
grid_data = pglib(”case39_epri”)
powerplot(case; layout_algorithm=Spring, iterations=50)
TABLE I: Time to compute layout [sec]
Layout Case39 Case118 Case500 Case1354 Case1888
Kamada Kawai 0.0096 0.12 2.3 31 35
Spring 0.0039 0.04 0.48 5 4.8
SFDP 0.00094 0.006 0.18 3.1 2.8
Spectral 0.0012 0.0085 0.097 2.5 2.2
Shell 0.00062 0.0042 0.013 0.068 0.065
Grid 0.00055 0.0022 0.015 0.25 0.065
Table I shows the computational time for each layout method on networks from 39 to 1,888 buses. On the largest network, the Kamada-Kawai layout requires 35 seconds to compute, while other methods may only require 2 to 4 seconds. The Shell and Grid layouts are trivial to calculate and require much less than one second to compute.
Figure 3: Partial fixed layout of power network. Bus data is known, but generator locations are not. The generator locations are computed using SFDP. The figure shows a plan for a Public Safety Power Shutoff plan to reduce wildfire risk by turning off power lines. Components in gray indicate that they have been de-energized [19].
A layout is not required if node location data is available for the network. If partial location data is available, then the the SFDP layout can compute locations for the unknown nodes. For example, Fig. 3 shows the RTS-GMLC system [38] which has location data for each bus, but the layout of generators is computed.
using PowerPlots, PGLib
grid_data = pglib(”case39_epri”)
# existing x and y coords are fixed
# missing coords are computed.
powerplot(grid_data; fixed=true)
A layout can also be precomputed and saved to the network data to avoid repeated computation for a large network.
using PowerPlots, PGLib
grid_data = pglib(”case39_epri”)
grid_data = layout_network(grid_data; layout_algorithm=SFDP, C=0.1, K=0.9)
powerplot(grid_data; fixed=true)
III-C Interactivity
A key value of PowerPlots.jl in exploring a power grid data file is the ability to interactively view the grid information. When hovering a mouse pointer over a power grid component in the plot, the associated information is shown, as seen in Fig. 4(b). Plotting a power grid can make it easy to visually identify patterns in the network, which can be further explored by hovering over the component to see valuable information. An example of the utility of this feature is shown in Fig. 4 where the locational marginal price (LMP) is plotted. The LMP in a network can vary significantly as a result of congestion. However, congestion is a difficult metric to gain intuition about solely from looking at numerical data. Fig. 4 shows the LMP for each node, and labels the congested branches in the network. The congestion pattern is quickly recognizable and a user can then hover over a node to see the power demand of a transmission line to find the parameter values of the line. This provides useful information to develop a deeper analysis of the network congestion.
((a))
((b))
Figure 4: Fig. 4(a) shows a plot of the Locational Marginal Price at each node, with binding transmission limits shown in red. In Fig. 4(b), a user hovers over a binding transmission line to see detailed information.
III-D Multi-Network Data
Multi-Network data is a data representation in PowerModels.jl for time series data, scenario data, multiple unique grid networks, or any other application where multiple representations of network data are utilized. The data file contains a new layer in the nested dictionary for each network in the multi-network data. The following code block shows an example of the multi-network data:
multinetwork_data = Dict(
nw => Dict(
”1” =>
bus => Dict(
”1” => Dict(
vmax => 1.1,
”2” => Dict( )
gen => Dict( )
”2” =>
bus => Dict(
”1” => Dict(
vmax => 1.1,
PowerPlots.jl uses this extra layer to create a filter in the plot to select which network is being viewed. This feature is useful, for example, in restoration planning to view the repair state of the network and the power flow at each stage of restoration. Fig. 5 shows an optimized restoration plan implemented from [16] to optimize the order of repairs of damaged power lines to minimize total un-served energy. The damaged branches in gray are sequentially repaired and turn green. In addition, the amount of power flow is indicated by the size of the branches. At repair step 24, two islands have been created, seen in Fig. 5(a). Fig. 5(b) shows repair 25, where a branch that connects these two islands is repaired and a significant amount of power flows from the left island to the right island. This indicates that the optimal solution for repair planning was to create multiple islands, before later connecting the islands.
((a)) Repair stage 24
((b)) Repair stage 25
Figure 5: Restoration plan applied to a 24-bus network. Repaired branches are shown in green, while damaged lines are shown in gray. The size of the branches indicates power flow. The restoration plan created two islands which are connected to each other late in the repair process. The left island contributes significant power flow to the top center nodes after they are connected.
III-E Supported Networks and Components
While originally designed for transmission networks, PowerPlots.jl also supports PowerModelsDistribution.jl [14] for distribution network visualization. PowerModelsDistribution.jl has two data models, an engineering model representing the actual components of the power grid, and a mathematical model representing an equivalent network that is used for its optimization problems. An example of the two network representations is shown in Fig. 6.
((a)) Engineering Model
((b)) Math Model
Figure 6: Distribution grids have a engineering model and a mathematical model. Shown here is the equivalent circuit of a 3-three center tap transformer.
This figure also shows the large number of component types that can be visualized. By default, all of these components are included in the figure. A subset of those components can be specified, and any additional components can also be specified. For example, if a researcher is creating a problem with a new component in the data model, e.g ”hydro”, they can plot the new component with
[…]
III-F Customization
Many of the features of PowerPlots.jl allow for quick visualization of a network to gain an understanding of the structure and patterns in the system. When communicating the information to an audience, a user will often reduce the information shown in the plot to create a figure tailored to the topic of their research. Because PowerPlots.jl uses the plotting library VegaLite, it is possible to take advantage of the many features in VegaLite to customize a final figure. Here, we take the LMP figure from Fig. 4 and show the changes from a basic plot to a polished figure. First, we create a plot with the following specification. The code to compute the LMP is not shown for brevity. As seen here, no connected components are shown, the size of the figure is set, and data fields of the bus and branch components are specified alongside the color scheme. These modifications are enabled by passing user arguments to PowerPlots.jl.
p = powerplot(data, connected_components=Symbol[],
height=250, width=250,
bus=(
:size=>100,
:data=>:LMP,
:data_type=>:quantitative,
:color=>reverse( colorscheme2array(
ColorSchemes.colorschemes[:RdYlGn_4]
branch=(
:size=>2,
:data=>:loading,
:data_type=>:nominal,
:color=>[”red”,”black”]
The figure created by this code is shown in Fig. 7. It is very similar to the image from Fig. 4(a) with some exceptions: 1) The legend titles are the component names, instead of the data they show. 2) The domain of the bus color range starts from $\approx 18$. These are changes that would need to be modified to accurately communicate results to a reader or viewer.
Figure 7: LMP figure with no customizations made of the output of the ‘powerplot()‘ function .
The following code edits the plot to make these changes and create the figure as shown in Fig. 4(a).
p.layer[1][”layer”][1][”encoding”][”color”] [”legend”] = Dict(”title”=>”p.layer[2][”encoding”][”color”][”scale”] [”domain”] = [0, max_LMP]
p.layer[2][”encoding”][”color”][”legend”] = Dict(”title”=>”LMP \\$/MWh”)
Additional modifications like modifying the size of each bus according to the power demand at the node, altering the height of the legend color gradient, and changing the font of the legend to match the manuscript is shown below, and the resulting figure shown in Fig. 8.
p.layer[2][”encoding”][”color”][”legend”] [”gradientLength”]=80
p.layer[2][”encoding”][”size”] = Dict{String,Any}(”field”=>”load”, type”=>”quantitative”)
p.layer[2][”encoding”][”size”][”legend”] = Dict(”title”=>”Load p.u.”)
p.layer[2][”encoding”][”size”][”scale”] = Dict(”range”=> [50,400])
[…]
@set! p.resolve.scale.size = :independent
[…]
for x in [”labelFont”, titleFont”]
p.layer[1][”layer”][1][”encoding”] [”color”][”legend”][x] = font
p.layer[2][”encoding”][”color”] [”legend”][x] = font
p.layer[2][”encoding”][”size”] [”legend”][x] = font
end
Figure 8: Power grid showing LMP by node color and nodal load by node size, with binding transmission limits in red. The legend titles have been modified, the limits of the color range for the LMP has been set to $0,\max\{LMP\}]$, and the height of the color range has been reduced. The font in the legend has been changed to match the font of this manuscript.
III-G Summary
The various features in this section highlight the many ways to use and interact with PowerPlots.jl. These features enable research by being easy to use for initial data exploration, allowing for flexibility in the data visualization, and supporting customized visualizations for publication.
IV Data Analysis
The data structures in PowerPlots.jl can also be used to enable data analysis of power grid data. This can provide insights into the network structure or patterns of data that otherwise can be difficult to identify.
IV-A Graph Analysis
[…]
Figure 9: Distribution of bus node-degree across all networks in PGLib. Small networks are less than 1,000 buses, medium between 1,000 and 10,000 buses, and large networks are greater than 10,000 buses.
IV-B Tabular Data Analysis
[…]
pmd = PowerModelsDataFrame(data)
sort!(pmd.components[:gen].pmax)
pmd.components[:gen][1:5,:index]
# the five highest capacity generators are
# 1, 54, 101, 41, and 65
Example: The following example explores metrics on the voltage magnitude at buses by voltage level. The code example compute the mean, standard deviation, minimum, and maximum voltage for each group of voltage levels.
data = pglib(”pglib_opf_case2000”)
result = solve_ac_opf(data, Ipopt.Optimizer)
update_data!(data, result[”solution”])
pmd = PowerModelsDataFrame(data)
gdf = groupby(pmd.components[:bus], :base_kv)
sort(combine(gdf, nrow, :vm => mean, :vm => std, :vm => minimum, :vm => maximum), :base_kv)
Table II shows these metrics. We find that the voltage is noticeable higher on the 69 kv buses compared to the high voltage network or to the low voltage network. This could be indicative of a significant amount of power injection occurring at that voltage level in this network.
Base kv Bus Count $|V|$ Mean $|V|$ Std $|V|$ Min $|V|$ Max
13 4 1.05 0.0414 1 1.1
14 121 1.04 0.0441 0.9 1.1
18 33 1.05 0.0324 0.933 1.1
20 14 1.04 0.045 0.935 1.09
22 13 1.04 0.0482 0.906 1.09
24 7 1.03 0.0521 0.934 1.08
69 686 1.07 0.0126 1.03 1.1
100 418 1.03 0.0523 0.911 1.1
138 526 1.06 0.0145 1.02 1.09
345 178 1.06 0.0319 0.929 1.09
TABLE II: Bus voltage magnitude metrics aggregated by bus voltage.
V Conclusion
Visualization of data is a powerful tool for building an understanding of data and highlighting aspects for communication. PowerPots.jl has a simple interface to construct a view of the network information, but has significant flexibility to enable visualization of any parameter of interest. It also allows significant customization to enable a user to create a figure that highlights the important information they wish to communicate. In creating PowerPlots.jl, several tools are developed that also provide utility in data analysis of the power grid. These include data transformations to graph network models and to tabular data formats that make applicable data analyses easier to perform. Together, the visualization and analysis tools in PowerPlots.jl enable researchers to explore novel power grid problems by providing capabilities that support research beyond standard power grid data. This software package will continue to evolve and develop new features. Additional customization features will be available from the user arguments, and the implementation of these features will be guided by community involvement in this package. In addition, future extensions of this work can be applied to visualizing other network systems such as gas, water, or telecommunications networks. This will eventually enable a visual understanding for cross-sector critical infrastructure analysis.
Acknowledgments
I would like to thank my PhD advisor Line Roald for her encouragement in developing this package, Hanbin Yang, Kshitij Girigoudar, and Joe Gorka for being early users of this package, and the people who have participated in the open source project by reporting issues and giving suggestions online. I would also like to thank the Advanced Network Science Initiative at Los Alamos National Laboratory for creating the InfrastructureModels.jl ecosystem on which PowerPlots.jl was developed. Support for this work was provided in part by the National Science Foundation (NSF) under the NSF CAREER award No. 2045860 and NSF ASCENT award No. 2132904, and by the Department of Energy through the LANL/LDRD Program and the Center for Nonlinear Studies.